OpenStudy (anonymous):

Owen made 100 sandwiches which she sold for exactly $100. She sold caviar sandwiches for $5.00 each, the bologna sandwiches for $2.00, and the liverwurst sandwiches for 10 cents. How many of each type of sandwich did she make?

OpenStudy (anonymous):

Tricky. Set up your equations: C = # of caviar B = # bologna L = # of liverwurst C + B + L = 100 5C + 2B + L/10 = 100 multiply the 2nd equation by 10 and subtract the two 49C + 19B = 900 Now, there are lots of ways to do that, but you need both C and B to be integers. Notice that 900 is divisible by 10. Divide the whole equation by 10 $4.9 C + $1.9 B = $90 Now, we know she need to make exactly $90 selling just caviar and bologna. Each time she sells a sandwich, she'll have to give a dime in change. To get back to an exact dollar figure, she'll need to sell some multiple of 10 sandwiches. So, make the substitution C + B = 10 n Now, solve for C and substitute 49 n - 3 B = 90 Remembering that n and B are both integers, n must be a multiple of 3 (divide that equation by 3 and see). So try, n=3, 6, 9, etc. For n=3, B = 19 For n=6, B = 68 For n=9, B = 117 (which is impossible.) If B = 19, then C = 11 and L = 70. If B = 68, then C = -8 which is impossible. So, the only possible answer is 11 caviar sandwiches, 19 bologna sandwiches and 70 Liverwurst sandwiches.

OpenStudy (hba):

Why post a question if you know the answer?

OpenStudy (anonymous):

Win.

OpenStudy (e.mccormick):

@hba Sometimes people want an answer checked.

OpenStudy (hba):

@e.mccormick Well that's something which is to be mentioned in the question.

OpenStudy (anonymous):

Yup

OpenStudy (e.mccormick):

True.